# On Arratia’s coupling and the Dirichlet law for the factors of a random integer

### v2

@@ Appendix A Tools from Number Theory
Understanding these functions gives information about the distribution of primes. In 1896, Charles de la Vallée Poussin and Jacques Hadamard proved the Prime Number Theorem, which gives an approximation for $\psi(x)$. The formulation we shall use in this paper is the following weaker version of their result:
@@
For a proof, see [12, Chapter 8].
Proposition A.2 (Strong Mertens’ estimate).
@@
By partial summation and the Prime Number Theorem (Proposition A.1), we directly have
$\sum_{p^{k}\leqslant x}\frac{1}{kp^{k}}=\int_{2^{-}}^{x}\frac{\mathrm{d}\psi(t)}{t\log t}=\log_{2}x+c+O\!\left(\frac{1}{(\log x)^{3}}\right)$
for some real constant $c$. Note that
$\sum_{p^{k}\leqslant x}\frac{1}{kp^{k}}+\log\prod_{p\leqslant x}\bigg{(}1-\frac{1}{p}\bigg{)}=-\sum_{\begin{subarray}{c}p^{k}>x\\ p\leqslant x\end{subarray}}\frac{1}{kp^{k}}.$
Since the right-hand side above tends to 00, and Mertens’ estimate says that
$\prod_{p\leqslant x}\bigg{(}1-\frac{1}{p}\bigg{)}\sim\frac{e^{-\gamma}}{\log x}$
(see [12, Theorem 3.4(c)] for a proof), then we have $c=\gamma$. ∎
@@
by the inequality $1+t\leqslant e^{t}$ and Mertens’ theorem (Proposition A.2). This completes the proof in this case. Finally, assume that $y\leqslant\sqrt{x}$. We then have
@@ B.1. The Total Variation Distance
(B.1) $d_{\mathrm{TV}}(\mu,\nu)\coloneqq\sup\left|\mu(A)-\nu(A)\right|,$
@@
$d_{\mathrm{TV}}(\mu,\nu)=\sum_{i\geqslant 1}(\mu(i)-\nu(i))^{+}=\sum_{i\geqslant 1}(\mu(i)-\nu(i))^{-}$
@@
Therefore, $d_{\mathrm{TV}}(\mu,\nu)=\sup_{A\subseteq\mathbb{N}}|\xi(A)|=\max\{\xi(E),-\xi(E^{c})\}$. Note that $\xi(E)=\sum_{i\geqslant 1}(\mu(i)-\nu(i))^{+}$, that $-\xi(E^{c})=\sum_{i\geqslant 1}(\mu(i)-\nu(i))^{-}$ and that
@@
  • $\mathbb{P}[X^{\prime}\neq Y^{\prime}]=d_{\mathrm{TV}}(X,Y)$.
We repeat his proof here. Let $\mu$ and $\nu$ be two probability measures supported on $\mathbb{N}$. Let $z_{j}\coloneqq\sum_{i\leqslant j}\frac{(\mu(i)-\nu(i))^{-}}{d_{\mathrm{TV}}(\mu,\nu)}$ (with $z_{0}\coloneqq 0$). We consider the function $f_{\mu,\nu}\colon\mathbb{N}\times(0,1)^{2}\to\mathbb{N}$ defined as
@@
Let $\mu$ and $\nu$ be two probability measures supported on $\mathbb{N}$, let $X$ be a random variable with law $\mu$, and let $U$ and $U^{\prime}$ be two uniform random variables in $(0,1)$ such that $X,U,U^{\prime}$ are independent. Let $Y\coloneqq f_{\mu,\nu}(X;U,U^{\prime})$ with $f_{\mu,\nu}$ defined as above. Then $\mathbb{P}[X\neq Y]=d_{\mathrm{TV}}(\mu,\nu)$ and $\mathbb{P}[Y\in A]=\nu(A)$.
@@
(B.5) $\mathbb{P}[X\neq Y]=\mathbb{P}[U\cdot\mu(X)>\nu(X)]=\sum_{m\geqslant 1}(\mu(m)-\nu(m))^{+}=d_{\mathrm{TV}}(\mu,\nu)$
@@
$=\Big{(}\mu(n)-\big{(}\mu(n)-\nu(n)\big{)}^{+}\Big{)}+d_{\mathrm{TV}}(\mu,\nu)\cdot\frac{(\mu(n)-\nu(n))^{-}}{d_{\mathrm{TV}}(\mu,\nu)}=\nu(n),$
where we used (B.3) and (B.5) to show that $\mathbb{P}\big{[}U\cdot\mu(X)>\nu(X)\big{]}=d_{\mathrm{TV}}(\mu,\nu)$. This concludes the proof of the lemma. ∎
@@ B.2. Poisson Processes
Appendix C Tools from Topology
Recall the following basic definitions.
Definition C.1.
Let $(X,\mathcal{O})$ be a topological space, and let $A\subseteq X$.
  1. (a)
    We define the interior of $A$ to equal $\text{int}(A)\coloneqq\{x\in X:\exists O\in\mathcal{O}\ \text{such that}\ x\in O\subseteq A\}$.
  2. (b)
    We define the closure of $A$ to equal $\overline{A}\coloneqq X\setminus\text{int}(X\setminus A)$.
  3. (c)
    We define the boundary of $A$ to equal $\partial A\coloneqq\overline{A}\setminus\text{int}(A)$.
  4. (d)
    We say that $A$ is disconnected if there exist two disjoint open sets $O_{1}$ and $O_{2}$ such that $A\subseteq O_{1}\sqcup O_{2}$ with $A\cap O_{1}$ and $A\cap O_{2}$ being both non-empty. We say that $A$ is connected if it is not disconnected.
Lemma C.2.
Let $X$ be a topological space and let $A,B\subseteq X$ such that $B\cap\partial A=\emptyset$ and $B$ is connected. Then either $B\subseteq A$ or $B\subseteq X\setminus A$.
Proof.
From the definition of closure and boundary above, the sets $\partial A,\text{int}(A)$ and $\text{int}(X\setminus A)$ form a partition of $X$. By our assumption that $B\cap\partial A=\emptyset$, we must thus have $B\subseteq\text{int}(A)\sqcup\text{int}(X\setminus A)$. By our assumption that $B$ is connected, we must then have that either $B\subseteq\text{int}(A)\subseteq A$ or that $B\subseteq\text{int}(X\setminus A)\subseteq X\setminus A$. This completes the proof. ∎

### v3

@@ Appendix A Tools from Number Theory
Understanding these functions gives information about the distribution of primes. In 1896, Charles de la Vallée Poussin and Jacques Hadamard proved the Prime Number Theorem, thus establishing that $\psi(x)\sim\theta(x)\sim x$ as $x\to\infty$. De la Vallée Poussin proved a more precise version of these estimates, with the error term $O\big{(}xe^{-c\sqrt{\log x}}\big{)}$ for some positive absolute constant $c$ (see, for example, [14, Chapter 8]). For the purposes of the present paper, however, the following weaker version of his result suffices:
@@
We shall also need a stronger version of Mertens’s estimate, which in its classical form, yields an estimate for the product $\prod_{p\leqslant x}(1-1/p)$.
Proposition A.2 (Strong Mertens’s estimate).
@@
Using the Prime Number Theorem (Proposition A.1), there exists a constant $c_{0}$ satisfying
$\sum_{p^{k}\leqslant x}\frac{1}{kp^{k}}=\int_{2^{-}}^{x}\frac{\,\mathrm{d}\psi(t)}{t\log t}=\log_{2}x+c_{0}+O\!\left(\frac{1}{(\log x)^{3}}\right).$
for $x\geqslant 2$. For the proof of $\gamma=c_{0}$, see the proof of [18, Theorem 2.7]. ∎
@@
by the inequality $1+t\leqslant e^{t}$ and Mertens’s theorem (Proposition A.2). This completes the proof in this case. Finally, assume that $y\leqslant\sqrt{x}$. We then have
@@ B.1. The Total Variation Distance
(B.1) $\mathrm{d}_{\mathrm{TV}}(\mu,\nu)\coloneqq\sup\left|\mu(A)-\nu(A)\right|,$
@@
$\mathrm{d}_{\mathrm{TV}}(\mu,\nu)=\sum_{i\geqslant 1}(\mu(i)-\nu(i))^{+}=\sum_{i\geqslant 1}(\mu(i)-\nu(i))^{-}$
@@
Therefore, $\mathrm{d}_{\mathrm{TV}}(\mu,\nu)=\sup_{A\subseteq\mathbb{N}}|\xi(A)|=\max\{\xi(E),-\xi(E^{c})\}$. Note that $\xi(E)=\sum_{i\geqslant 1}(\mu(i)-\nu(i))^{+}$, that $-\xi(E^{c})=\sum_{i\geqslant 1}(\mu(i)-\nu(i))^{-}$ and that
@@
  • $\mathbb{P}[X^{\prime}\neq Y^{\prime}]=\mathrm{d}_{\mathrm{TV}}(X,Y)$.
We repeat his proof here. Let $\mu$ and $\nu$ be two probability measures supported on $\mathbb{N}$. Let $z_{j}\coloneqq\sum_{i\leqslant j}\frac{(\mu(i)-\nu(i))^{-}}{\mathrm{d}_{\mathrm{TV}}(\mu,\nu)}$ (with $z_{0}\coloneqq 0$). We consider the function $f_{\mu,\nu}\colon\mathbb{N}\times(0,1)^{2}\to\mathbb{N}$ defined as
@@
Let $\mu$ and $\nu$ be two probability measures supported on $\mathbb{N}$, let $X$ be a random variable with law $\mu$, and let $U$ and $U^{\prime}$ be two uniform random variables in $(0,1)$ such that $X,U,U^{\prime}$ are independent. Let $Y\coloneqq f_{\mu,\nu}(X;U,U^{\prime})$ with $f_{\mu,\nu}$ defined as above. Then $\mathbb{P}[X\neq Y]=\mathrm{d}_{\mathrm{TV}}(\mu,\nu)$ and $\mathbb{P}[Y\in A]=\nu(A)$.
@@
(B.5) $\mathbb{P}[X\neq Y]=\mathbb{P}[U\cdot\mu(X)>\nu(X)]=\sum_{m\geqslant 1}(\mu(m)-\nu(m))^{+}=\mathrm{d}_{\mathrm{TV}}(\mu,\nu)$
@@
$=\Big{(}\mu(n)-\big{(}\mu(n)-\nu(n)\big{)}^{+}\Big{)}+\mathrm{d}_{\mathrm{TV}}(\mu,\nu)\cdot\frac{(\mu(n)-\nu(n))^{-}}{\mathrm{d}_{\mathrm{TV}}(\mu,\nu)}=\nu(n),$
where we used (B.3) and (B.5) to show that $\mathbb{P}\big{[}U\cdot\mu(X)>\nu(X)\big{]}=\mathrm{d}_{\mathrm{TV}}(\mu,\nu)$. This concludes the proof of the lemma. ∎